OpenStudy (anonymous):

Are the events, earn an A and study, independent or dependent? What is the probability of randomly selecting a student who did not earn an A given that the student studied? Express the answer as a fraction in simplest form

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (amistre64):

if the probability of earning an A is the same for all given event that A occurs in, then the probability of A is not dependant on any particular event.

OpenStudy (amistre64):

what is the P(A|study)? what is the P(A|total events) ? are they the same? or does the probability of A changed depending on which event takes place?

OpenStudy (amistre64):

is 6 out of 30, proportionate to 7 out of 34?

OpenStudy (anonymous):

yes because some students who did study got an A

OpenStudy (amistre64):

thats not whats being asked.

OpenStudy (amistre64):

the probability of A given then event'study': 6 out of 30 right?

OpenStudy (anonymous):

yes correct

OpenStudy (amistre64):

the probability of A given the totality of all events: 7 out of 34 right?

OpenStudy (anonymous):

yes correct

OpenStudy (amistre64):

since the probability of A, depends on the event in which it occurs: 6/30 not= 7/34 then A depends on a particular event. They are not independant.

OpenStudy (amistre64):

agreed?

OpenStudy (anonymous):

yes i do agree

OpenStudy (amistre64):

good, just making sure we are not confusing "mutually exclusive" with "independence" so, do we know how to calculate what they are asking for now?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it owuld be 6/30 and 7/34

OpenStudy (amistre64):

read the question carefully the probability of NOTA, given that they studied. A, given study is 6/30 .... 1 - A,given study is NOTA, given study.

OpenStudy (amistre64):

this is just a depressing question, it ultimately shows that the majority of students studied and still failed ....

OpenStudy (amistre64):

88% of the kids studied, and of them, 80% failed to get an A ....

OpenStudy (anonymous):

It reflects negatively on the teacher :)

OpenStudy (amistre64):

the teachers i had that graded on a curve used that philosophy. They assumed if the majority of the students did not get a good grade, it was because they had not been tested at their skill level - they had not learned the material.

OpenStudy (anonymous):

lol dont you agree though??? Or do you think ppl dont like to stretch their brains and think out of the box? I really do not think its pure laziness ...

OpenStudy (amistre64):

i most likely irritated my science class. I was taking the tests and all the correct answer were darker ink then the others. After about 4 questions i brought it to his attention. He had mistakenly printed out the answer key instead. He had to collect all the tests and randomize the question pool again.

OpenStudy (amistre64):

IF they studied, then their performance is caused by something else. IF they fail to study, its their own fault then

OpenStudy (anonymous):

so would the answer be B

OpenStudy (anonymous):

Lmaoooo Dude yea i ok i agree with that

OpenStudy (amistre64):

24 out of 30 24/30 does nto simplify to B

OpenStudy (amistre64):

6/30 = 1/5 24/30 = 4/5

OpenStudy (anonymous):

its 4/5 so its dependent correct??

OpenStudy (amistre64):

correct

OpenStudy (anonymous):

so the answer is A?

OpenStudy (amistre64):

thats the one id go with, but im not grading it